Hitting the hammer with the nail lasted delta t = 0.05 s. Calculate the impact force that acted on the nail, when it is known that the impulse of the hammer changed with delta p = 20 kgm / s.


1
Expert's answer
2021-01-31T10:23:07-0500
"F\Delta t=\Delta p,""F=\dfrac{\Delta p}{\Delta t}=\dfrac{20\ \dfrac{kgm}{s}}{0.05\ s}=400\ N."